OpenStudy (openstudygirl1):

Express 8.21 as a mixed number. Your number should be, a mixed number like 13/4. Will fan and medal Please help Me.

OpenStudy (adrimit):

the decimal terminates, how can i turn 8.32 into an integer?

OpenStudy (openstudygirl1):

ok and?

OpenStudy (adrimit):

sorry 8.21

OpenStudy (openstudygirl1):

i dont understand

OpenStudy (adrimit):

what happens when i multiply 8.21 by 10?

OpenStudy (openstudygirl1):

82.1

OpenStudy (adrimit):

ok do it again...

OpenStudy (openstudygirl1):

82.1 ??????

OpenStudy (adrimit):

no i mean what is 10 times 82.1?

OpenStudy (openstudygirl1):

ohh

OpenStudy (openstudygirl1):

811

OpenStudy (openstudygirl1):

is the answer 8/11 ?

OpenStudy (adrimit):

you're overthinking it

OpenStudy (openstudygirl1):

?

OpenStudy (adrimit):

just move the decimal like you did before

OpenStudy (openstudygirl1):

ummm...

OpenStudy (openstudygirl1):

how???????

OpenStudy (adrimit):

ok, when you multiply a decimal number by 10, the position of the decimal point moves one place to the right

OpenStudy (adrimit):

that is why 8.21 * 10 = 82.1

OpenStudy (openstudygirl1):

whats the Answer then?

OpenStudy (adrimit):

i am trying to teach you in a way that you can do any problem of this type on your own. I am basically walking you through my thought process to get the answer.

OpenStudy (openstudygirl1):

ok

OpenStudy (openstudygirl1):

i almost have my break soon

OpenStudy (adrimit):

ok i'll go faster.

OpenStudy (openstudygirl1):

thank you

OpenStudy (adrimit):

8.21 * 10^2 = 821 (move the decimal two places to the right) 8.21=821/100 this is the final answer since the numerator and denominator have no prime factors in common

OpenStudy (openstudygirl1):

ok

OpenStudy (adrimit):

the method is to multiply by 10^n where n=number of decimal places (only if finite)

OpenStudy (openstudygirl1):

so the answer would be then 821/100?

OpenStudy (adrimit):

exactly

OpenStudy (openstudygirl1):

Thank you so much

jimthompson5910 (jim_thompson5910):

\(\Large 8.{\color{red}{21}}\) means \(\Large 8 \text{ and }{\color{red}{21}} \text{ hundredths }\) Since the "1" in \(\Large {\color{red}{21}}\) is in the hundredths place. So you simply place that 21 over 100 like so \(\Large 8.{\color{red}{21}} = 8 + \frac{{\color{red}{21}}}{100}\) \(\Large 8.{\color{red}{21}} = 8\frac{{\color{red}{21}}}{100}\)

OpenStudy (openstudygirl1):

is that it?

OpenStudy (adrimit):

that works too

jimthompson5910 (jim_thompson5910):

Another example: \[\Large 7.219 = 7\frac{219}{1000}\] the "9" in "219" is in the thousandths decimal place, so we place 219 over 1000

OpenStudy (adrimit):

what about 9.40404040...?

jimthompson5910 (jim_thompson5910):

Let x = 9.40404040... Multiply both sides by 100 to get 100x = 940.404040... Now subtract 100x - x = 940.404040... - 9.40404040... 99x = 931 Then solve for x 99x = 931 x = 931/99 So 931/99 = 9.404040404040...

jimthompson5910 (jim_thompson5910):

the improper fraction \(\Large \frac{931}{99}\) is the same as the mixed number \(\Large 9\frac{40}{99}\)